# Graph Neural Networks for Next-Generation-IoT: Recent Advances and Open Challenges

### v2

@@ Abstract
Graph Neural Networks (GNNs) have emerged as a powerful framework for modeling complex interconnected systems, hence making them particularly well-suited to address the growing challenges of next-generation Internet of Things (NG-IoT) networks. Despite increasing interest in this area, existing studies remain fragmented, and there is a lack of comprehensive guidance on how GNNs can be systematically applied to NG-IoT systems. As NG-IoT systems evolve toward 6G, they incorporate diverse technologies such as massive MIMO, reconfigurable intelligent surfaces (RIS), terahertz (THz) communication, satellite systems, mobile edge computing (MEC), and ultra-reliable low-latency communication (URLLC). These advances promise unprecedented connectivity, sensing, and automation but also introduce significant complexity, requiring new approaches for scalable learning, dynamic optimization, and secure, decentralized decision-making. This survey provides a comprehensive and forward-looking exploration of how GNNs can empower NG-IoT environments structured as ten open research questions that span the relevant theoretical foundations, practical deployments, and emerging integration pathways. We commence by exploring the fundamental paradigms of GNNs and articulating the motivation for their use in NG-IoT networks. Besides, to further justify their suitability, we intrinsically connect GNNs for the first time with the family of low-density parity-check (LDPC) codes, modeling the NG-IoT as dynamic constrainted graphs where GNNs harness belief propagation for convergence and interpretability through density evolution and EXIT charts. We highlight the distinct roles of node-, edge-, and graph-level tasks in tackling key challenges and demonstrate the GNNs’ ability to overcome the limitations of traditional optimization methods. Following this, we examine the application of GNNs across core NG-enabling technologies and their integration with distributed frameworks to support privacy-preservation and distributed intelligence. We then delve into the challenges posed by adversarial attacks, offering insights into defense mechanisms to secure GNN-based NG-IoT networks. Lastly, we examine how GNNs can be integrated with emerging technologies like integrated sensing and communication (ISAC), satellite-air-ground-sea integrated networks (SAGSIN), and quantum computing. Our findings highlight the transformative potential of GNNs in improving efficiency, scalability, and security within NG-IoT systems, paving the way for future advances. Finally, we summarize the key lessons learned throughout the paper and outline promissing future research directions, along with a set of design guidelines aimed at facilitating the development of efficient, scalable, and secure GNN models tailored for NG-IoT applications.
@@ I Introduction
1) What Are Graphs and Graph Neural Networks? Graph definition and Graph Neural Networks (GNNs) paradism Graph data, GNN Paradigm, and GNN variants [29, 30, 31] The explainability and scalability of GNNs [32, 33] Model depth and over-smoothing Scalability of GNN models Security
2) Why GNNs for NG-IoT Networks? How Do We Model NG-IoT Systems as Graphs? NG-IoT networks feature dynamic topology, heterogeneity, and non-Euclidean structures, making them well-suited for graph modeling. GNNs enable effective learning at node, edge, and graph levels to address these complexities. Node-level task [34, 35, 36, 37, 38, 39, 40, 41, 42] Edge-level task [43, 44, 45, 46, 47] Graph-level task [48, 49, 50, 51] High computational and memory requirements. Divide large graphs into subgraphs for local learning and then aggregate Adapting DE and EXIT chart for analyze GNN and model design
@@ II-1 Background:
A graph is a mathematical representation consisting of a set of vertices (also called nodes) and edges (also called links) that connect pairs of vertices. In general, a graph is represented by a tuple $\mathcal{G}=(\mathcal{V},\mathcal{E})$. The set of $|\mathcal{V}|$ vertices is denoted by $\mathcal{V}=\{1,...,v,...|\mathcal{V}|\}$, while the set of edges is defined based on the graph types: Undirected graph: =$\mathcal{E}=\{(i,j)|i,j\in\mathcal{V},i\neq j\}$, where $(i,j)$ represents an undirected connection between vertices $i$ and $j$; Directed graph: =$\mathcal{E}=\{(i,j)|i,j\in\mathcal{V},i\neq j\}$, where $(i,j)$ is the edge directed from vertex $i$ to vertex $j$. Graphs can also be categorized into two types based on the nature of their nodes and edges: Homogeneous graph: Consists of a single type of node and a single type of edge; Heterogeneous graph: These graphs include multiple types of nodes and edges. The illustrations of different types of graphs are represented in Fig. 2. This diversity better represents complex real-world networks, such as wireless communication systems, where different devices and connections exist. The adjacency matrix of the graph is represented as $\mathbf{A}\in\{0;1\}^{|\mathcal{V}|\times|\mathcal{V}|}$, where $\mathbf{A}_{ij}=1$ if ${(i,j)}\in\mathcal{E}$. For an undirected graph, $\mathbf{A}$ is symmetric, while for a directed graph, it may not be. The node feature matrix $\mathbf{X}$ is a $|\mathcal{V}|\times F_{n}$ matrix, where each row corresponds to a vector of features for a node. Similarly, the edge feature matrix $\mathbf{E}$ is a ||×$|\mathcal{E}|\times F_{e}$ matrix, where each row corresponds to features of an edge. By storing information in the nodes and edges, the graph can capture the complexity of real-world networks, as illustrated in Fig. 3. In Fig. 3(a), a hierarchical wireless communication network is represented as a graph, where each system entity, including user equipment (UEs), unmanned aerial vehicles (UAVs), and access points (APs), is modeled as a node. Communication links between entities are depicted as edges, with different edge types capturing distinct connection types. For example, the edge between an AP and the $i$-th UAV is denoted as $e^{\textrm{U-A}}{1}=[g_{i}]$, while the edge between the $i$-th UAV and the $j$-th UE is denoted as $e^{\textrm{U-U}}{ij}=[h_{ij}]$, representing their respective channel gains. Fig. 3(b) illustrates another graph-based example: a chemical structure. Here, atoms such as Carbon (C), Oxygen (O), Nitrogen (N), and Sulfur (S) are represented as nodes, while chemical bonds (single, double, or aromatic) are represented as undirected edges. Fig. 3(c) shows graph modeling in social networks, where individuals are nodes and their interactions, such as friendships, follows, or message exchanges, form the edges. These edges can be directed or undirected, depending on the nature of the relationship. For instance, a “follow” on platforms like Instagram or Twitter is a directed edge, while a mutual friendship results in an undirected edge. Such graph representations are widely used in applications like recommendation systems, influence propagation, and community detection.
Figure 4: A general pipeline for solving wireless network problems by using GNNs.
@@
  • How GNNs works?: As the GNN processes the graph, each node’s embedding vector is iteratively updated by aggregating its own features along with information from neighboring nodes and edges, as shown in the graph convolutional layer block in Fig. 4. In contrast, MLP treats each node independently without leveraging structural relationships, aggregating information only from fully connected layers, while CNN applies local filters to structured grid-based data, capturing spatial dependencies within a fixed neighborhood but lacking adaptability to irregular graph structures. Ultimately, each node and edge obtains an embedding that captures the broader network context. The GNN model then processes these node and edge representations to produce node-level, edge-level, or graph-level features for tasks such as node clustering, link prediction, or graph classification. Fig. 4 illustrates the general pipeline of a GNN model, comprising input data, graph convolutional layers, and a downstream task layer.
@@ II-2 State-of-the-art:
  • Node level: Node-level GNN applications are widely used in wireless networks, especially in power allocation [34, 35, 39]. For instance, in [20], the uplink power allocation of a cell-free massive MIMO IoT System can be formulated as a node-level task of an undirected graph, where users and access points (APs) are represented as nodes. This approach allows the model to generalize effectively across different network configurations, ensuring scalability. However, in downlink power allocation, the problem becomes more complex because each AP must allocate power vectors that are dependent on the number of served users, making the task more sensitive to changes in network size. To overcome this challenge, Shen et al. [152] proposes to represent communication links as nodes, as illustrated in Fig. 7. This representation shifts the focus to the links, allowing the system to maintain scalability even as the network size varies. This method ensures that the power allocation process remains efficient and adaptable in dynamic network environments, where the number of connected devices may fluctuate. Furthermore, by representing each communication link as a node, the node-level tasks lend themselves to solving link scheduling problems, such as AP-IoT device association or frequency assignment problems. In this way, decisions regarding transceiver pair scheduling can be made independently based on the node’s features. This method enables efficient link scheduling, as demonstrated in studies like [36, 37].
@@
    The results show that an edge-GNN and node-GNN constitute a pair of models capable of approximating the optimal solution. Briefly, these two models achieve a minimum rate above $1$bps/Hz in about $65$% of the instances. The hybrid quantum GNN, which integrates quantum circuits into the Node-GNN architecture, also performs competitively. Quantitatively, this model tends to achieve $1$bps/Hz in about $64$% of the instances. On average, the Edge-GNN, Node-GNN, and HQGNN models provide data throughputs of $1.0783$, $1.0733$, and $1.0640$ bps/Hz, respectively. By contrast, the optimal solution achieves an average throughput of $1.129$ bps/Hz, slightly outperforming all GNN models, but at the cost of significantly higher computational complexity, particularly as the network size escalates. The graph-GNN performs worst of all GNN models, with less than $40$% of cases reaching a minimum rate of $1.0$ bps/Hz. This poor performance is attributed to its global aggregation of information, which fails to capture the crucial local node and edge interactions necessary for optimizing power allocation in the network. Overall, the results underscore the superiority of node-level GNN models, particularly of those utilizing attention mechanisms for solving the power control problem of cell-free massive MIMO systems. Furthermore, the hybrid quantum GNN offers an intriguing avenue for integrating quantum computing into GNN architectures.
@@ II-2 State-of-the-arts
Model Optimization for Edge and Cloud Deployment: To streamline the deployment of GNNs in NG-IoT networks, various studies have explored different key aspects. In attempt to embed GNN, Zhang et al. [52] have explored how (FPGAs) can accelerate message-passing operations in GNNs, leading to faster inference times while maintaining model accuracy. Besides, the authors also summarizes the challenges that FPGA-based GNNs accelerators need to address. Meanwhile, Wang et al. [53] proposes a quantization-based approach to address scalability challenges in GNNs. By compressing models with low-bit quantization and controlling message propagation, it reduces computational and memory costs while maintaining accuracy. The method achieves significant speedups (5.11×5.$1$ for INT2, 4.70×4.70$\times$4.70 × for INT4), offering a cost-effective solution for deploying GNNs in NG-IoT networks. Xie et al. [54] proposed Graph sparsification and network pruning (GASSIP), a novel framework for lightweight graph neural architecture search (GNAS) tailored for resource-constrained scenarios. By combining operation-pruned architecture search and curriculum graph sparsification, GASSIP efficiently identifies optimal lightweight GNNs with reduced model parameters and sparser graphs. Extensive experiments show that GASSIP achieves comparable or superior performance with significantly fewer resources, making it highly relevant for cost-effective GNN deployment in NG-IoT applications. In [55], Wu et al. introduced Prototype-Guided Knowledge Distillation (PGKD), a novel method for distilling high-accuracy GNNs into low-latency MLPs without relying on graph edges. By leveraging class prototypes and structure-aware distillation losses, PGKD effectively transfers graph structural knowledge from GNNs to MLPs, enabling edge-free learning.
Multi-modal data processing in NG-IoT networks: Li et al. [56] proposed GNNMR, a multi-modal recommendation framework that integrates GNNs with deep mutual learning to address modality bias. By training separate GNNs on uni-modal user-item graphs and using mutual knowledge distillation, GNNMR synchronizes latent semantic relationships across modalities, improving multi-modal embeddings. It outperforms existing methods in Top-K recommendation tasks, making it suitable for NG-IoT applications requiring efficient user-item interaction modeling. Similarly, Zhang et al. [57] introduced a GNN framework for multimodal data integration, leveraging feature, decision, and deep fusion techniques to achieve high accuracy ($98$%), recall ($86.9$%), and F1 score ($0.964$). This approach demonstrates robust performance in complex scenarios, highlighting the potential of multimodal GNNs for NG-IoT applications.

### v3

@@ Abstract
Graph Neural Networks (GNNs) have emerged as a powerful framework for modeling complex interconnected systems, hence making them particularly well-suited to address the growing challenges of next-generation Internet of Things (NG-IoT) networks. Despite increasing interest in this area, existing studies remain fragmented, and there is a lack of comprehensive guidance on how GNNs can be systematically applied to NG-IoT systems. As NG-IoT systems evolve toward 6G, they incorporate diverse technologies such as massive MIMO, reconfigurable intelligent surfaces (RIS), terahertz (THz) communication, satellite systems, mobile edge computing (MEC), and ultra-reliable low-latency communication (URLLC). These advances promise unprecedented connectivity, sensing, and automation but also introduce significant complexity, requiring new approaches for scalable learning, dynamic optimization, and secure, decentralized decision-making. This survey provides a comprehensive and forward-looking exploration of how GNNs can empower NG-IoT environments structured as ten open research questions that span the relevant theoretical foundations, practical deployments, and emerging integration pathways. We commence by exploring the fundamental paradigms of GNNs and articulating the motivation for their use in NG-IoT networks. Besides, to further justify their suitability, we intrinsically connect GNNs for the first time with the family of low-density parity-check (LDPC) codes, modeling the NG-IoT as dynamic constrainted graphs where GNNs harness belief propagation for convergence and interpretability through density evolution and EXIT charts. We highlight the distinct roles of node-, edge-, and graph-level tasks in tackling key challenges and demonstrate the GNNs’ ability to overcome the limitations of traditional optimization methods. Following this, we examine the application of GNNs across core NG-enabling technologies and their integration with distributed frameworks to support privacy-preservation and distributed intelligence. We then delve into the challenges posed by adversarial attacks, offering insights into defense mechanisms to secure GNN-based NG-IoT networks. Lastly, we examine how GNNs can be integrated with emerging technologies like integrated sensing and communication (ISAC), satellite-air-ground-sea integrated networks (SAGSIN), and quantum computing. Our findings highlight the transformative potential of GNNs in improving efficiency, scalability, and security within NG-IoT systems, paving the way for future advances. Finally, we summarize the key lessons learned throughout the paper and outline promising future research directions, along with a set of design guidelines aimed at facilitating the development of efficient, scalable, and secure GNN models tailored for NG-IoT applications.
@@ I Introduction
2023 GraphGPT Combined large language models(LLMs) with graph structural learning via instruction tuning, enabling cross-domain generalization Generalized graph representation across datasets with zero-shot and supervised learning using text-graph alignment (Zero-Shot Learning, Instruction-Tuned LLMs, Cross-Modal Graph Learning) Tang, J. et al. [13]
2024 Retrieval-Augmented GNNs Unified retrieval + GNN framework to inject non-parametric knowledge into graph learning, improving generalization for unseen entities or relations Knowledge Graphs, Open-World QA, Recommender Systems Han, H. et al. [14]
@@
1) What Are Graphs and Graph Neural Networks? Graph definition and Graph Neural Networks (GNNs) paradism Graph data, GNN Paradigm, and GNN variants [31, 32, 33] The explainability and scalability of GNNs [34, 35] Model depth and over-smoothing Scalability of GNN models Security Privacy-specific to GNN
2) Why GNNs for NG-IoT Networks? How Do We Model NG-IoT Systems as Graphs? NG-IoT networks feature dynamic topology, heterogeneity, and non-Euclidean structures, making them well-suited for graph modeling. GNNs enable effective learning at node, edge, and graph levels to address these complexities. Node-level task [36, 37, 38, 39, 40, 41, 42, 43, 44, 45] Edge-level task [46, 47, 48, 49, 50] Graph-level task [51, 52, 53, 54] High computational and memory requirements. Divide large graphs into subgraphs for local learning and then aggregate Adapting DE and EXIT chart for analyze GNN and model design
@@ II-1 Background:
A graph is a mathematical representation consisting of a set of vertices (also called nodes) and edges (also called links) that connect pairs of vertices. In general, a graph is represented by a tuple $\mathcal{G}=(\mathcal{V},\mathcal{E})$. The set of $|\mathcal{V}|$ vertices is denoted by $\mathcal{V}=\{1,...,v,...|\mathcal{V}|\}$, while the set of edges is defined based on the graph types: Undirected graph: $\mathcal{E}=\{(i,j)|i,j\in\mathcal{V},i\neq j\}$, where $(i,j)$ represents an undirected connection between vertices $i$ and $j$; Directed graph: $\mathcal{E}=\{(i,j)|i,j\in\mathcal{V},i\neq j\}$, where $(i,j)$ is the edge directed from vertex $i$ to vertex $j$. Graphs can also be categorized into two types based on the nature of their nodes and edges: Homogeneous graph: Consists of a single type of node and a single type of edge; Heterogeneous graph: These graphs include multiple types of nodes and edges. The illustrations of different types of graphs are represented in Fig. 2. This diversity better represents complex real-world networks, such as wireless communication systems, where different devices and connections exist. The adjacency matrix of the graph is represented as $\mathbf{A}\in\{0;1\}^{|\mathcal{V}|\times|\mathcal{V}|}$, where $\mathbf{A}_{ij}=1$ if ${(i,j)}\in\mathcal{E}$. For an undirected graph, $\mathbf{A}$ is symmetric, while for a directed graph, it may not be. The node feature matrix $\mathbf{X}$ is a $|\mathcal{V}|\times F_{n}$ matrix, where each row corresponds to a vector of features for a node. Similarly, the edge feature matrix $\mathbf{E}$ is a $|\mathcal{E}|\times F_{e}$ matrix, where each row corresponds to features of an edge. By storing information in the nodes and edges, the graph can capture the complexity of real-world networks, as illustrated in Fig. 3. In Fig. 3(a), a hierarchical wireless communication network is represented as a graph, where each system entity, including user equipment (UEs), unmanned aerial vehicles (UAVs), and access points (APs), is modeled as a node. Communication links between entities are depicted as edges, with different edge types capturing distinct connection types. For example, the edge between an AP and the $i$-th UAV is denoted as $e^{\textrm{U-A}}{1}=[g_{i}]$, while the edge between the $i$-th UAV and the $j$-th UE is denoted as $e^{\textrm{U-U}}{ij}=[h_{ij}]$, representing their respective channel gains. Fig. 3(b) illustrates another graph-based example: a chemical structure. Here, atoms such as Carbon (C), Oxygen (O), Nitrogen (N), and Sulfur (S) are represented as nodes, while chemical bonds (single, double, or aromatic) are represented as undirected edges. Fig. 3(c) shows graph modeling in social networks, where individuals are nodes and their interactions, such as friendships, follows, or message exchanges, form the edges. These edges can be directed or undirected, depending on the nature of the relationship. For instance, a “follow” on platforms like Instagram or Twitter is a directed edge, while a mutual friendship results in an undirected edge. Such graph representations are widely used in applications like recommendation systems, influence propagation, and community detection.
Figure 4: A general pipeline for solving wireless network problems using GNNs. The pipeline illustrates how graph-structured data, such as users, base stations, and links, is processed via graph convolutional layers to generate embeddings that facilitate downstream tasks like user association, resource allocation, and interference management.
@@
  • How GNNs work?: As the GNN processes the graph, each node’s embedding vector is iteratively updated by aggregating its own features with information gleaned from neighboring nodes and edges, as shown in the graph convolutional layer block of Fig. 4. In the context of wireless networks, this framework allows the modeling of complex relationships of users as nodes and their links (e.g., interference, connectivity, or cooperation) as edges. GNNs can learn meaningful representations that capture spatial-, temporal-, and resource-related dependencies, thereby enabling efficient solutions for tasks such as user association, frequency and power allocation, relay selection, and mobility prediction. The final node or edge embeddings are then input to the downstream layer in support of task-specific decisions. In contrast, MLP treats each node independently without leveraging structural relationships, aggregating information only from fully connected layers, while CNN applies local filters to structured grid-based data, capturing spatial dependencies within a fixed neighborhood but lacking adaptability to irregular graph structures. Ultimately, each node and edge obtains an embedding that captures the broader network context. The GNN model then processes these node and edge representations to produce node-level, edge-level, or graph-level features for tasks such as node clustering, link prediction, or graph classification. Fig. 4 illustrates the general pipeline of a GNN model, comprising input data, graph convolutional layers, and a downstream task layer.
@@
Privacy-specific to GNNs: The relational learning paradigm of GNNs introduces distinct privacy challenges that are fundamentally different from those in traditional deep learning architectures. In contrast to conventional neural networks that process isolated data points, GNNs inherently expose sensitive information through their graph-structured computations, hence resulting in vulnerabilities. Structural privacy risks arise from the inherent nature of message passing in GNNs, where iterative aggregation may inadvertently expose critical topological features, such as bridge links in social networks, proprietary dependencies in industrial IoT, or cluster memberships in critical infrastructure, potentially revealing sensitive relationships and operational structures [141, 142]. In these scenarios, protecting such high-impact connections is far more important than safeguarding non-essential links, whose disclosure typically has negligible operational impact. Attribute inference further aggravates these risks. The inference of embeddings and/or gradients can be exploited by adversaries for reconstructing sensitive node or edge attributes, posing significant privacy threats. Indeed, recent studies show that node attributes can be reconstructed from gradient updates [143] or final embeddings, while edge properties, such as weights and connection types, may be deduced by observing multiple GNN layers [144, 145].
To mitigate these risks, advanced privacy-preserving techniques have been proposed. For example, Yuan et al. [146] introduced locally private GNN training frameworks by allocating privacy budgets based on node degrees and aggregation perturbation methods to enforce personalized differential privacy (DP) during message passing, reducing leakage while maintaining utility. Furthermore, personalized DP mechanisms allow heterogeneous privacy budgets, enabling fine-grained protection across different nodes and edges. As a further advance, Li et al. [145] conceived node-level privacy-preserving embeddings for protecting sensitive attributes, while Sun et al. [147] developed decentralized DP approaches for the estimation of subgraph statistics, ensuring the protection of critical graph connections without sacrificing accuracy. Incorporating these state-of-the-art DP solutions represents a promising direction for enhancing privacy in GNN-based NG-IoT networks.
@@ II-3 Challenges and future directions:
  • Privacy: While adversarial security has received growing attention, privacy risks specific to GNNs remain underexplored. The message-passing mechanisms of GNNs can inadvertently expose sensitive topological or contextual information, such as bridge links or critical community structures, even in the absence of attacks [141]. Generic anonymization or link-hiding methods may protect non-essential edges but fail to safeguard high-impact relationships. Addressing this challenge requires research into context-aware privacy-preserving methods that specifically target critical connections and sensitive attributes in graph data.
@@ II-2 State-of-the-art:
L. Giang et al. [38] (2024) Wireless Sensor Networks Power control Undirected + Heterogeneous
@@
  • Node level: Node-level GNN applications are widely used in wireless networks, especially in power allocation [36, 37, 38, 42]. For instance, in [22], the uplink power allocation of a cell-free massive MIMO IoT System can be formulated as a node-level task of an undirected graph, where users and access points (APs) are represented as nodes. This approach allows the model to generalize effectively across different network configurations, ensuring scalability. However, in downlink power allocation, the problem becomes more complex because each AP must allocate power vectors that are dependent on the number of served users, making the task more sensitive to changes in network size. To overcome this challenge, Shen et al. [162] proposes to represent communication links as nodes, as illustrated in Fig. 7. This representation shifts the focus to the links, allowing the system to maintain scalability even as the network size varies. This method ensures that the power allocation process remains efficient and adaptable in dynamic network environments, where the number of connected devices may fluctuate. Furthermore, by representing each communication link as a node, the node-level tasks lend themselves to solving link scheduling problems, such as AP-IoT device association or frequency assignment problems. In this way, decisions regarding transceiver pair scheduling can be made independently based on the node’s features. This method enables efficient link scheduling, as demonstrated in studies like [39, 40].
@@
    The results show that an edge-GNN and node-GNN constitute a pair of models capable of approximating the optimal solution. Briefly, these two models achieve a minimum rate above 11bps/Hz in about 6565% of the instances. The hybrid quantum GNN, which integrates quantum circuits into the Node-GNN architecture, also performs competitively. Quantitatively, this model tends to achieve 11bps/Hz in about 6464% of the instances. On average, the Edge-GNN, Node-GNN, and HQGNN models provide data throughputs of $1.0783$, $1.0733$, and $1.0640$ bps/Hz, respectively. By contrast, the optimal solution achieves an average throughput of $1.129$ bps/Hz, slightly outperforming all GNN models, but at the cost of significantly higher computational complexity, particularly as the network size escalates. The graph-GNN performs worst of all GNN models, with less than 4040% of cases reaching a minimum rate of $1.0$ bps/Hz. This poor performance is attributed to its global aggregation of information, which fails to capture the crucial local node and edge interactions necessary for optimizing power allocation in the network. Overall, the results underscore the superiority of node-level GNN models, particularly of those utilizing attention mechanisms for solving the power control problem of cell-free massive MIMO systems. Furthermore, the hybrid quantum GNN offers an intriguing avenue for integrating quantum computing into GNN architectures.
@@ II-2 State-of-the-arts
Model Optimization for Edge and Cloud Deployment: To streamline the deployment of GNNs in NG-IoT networks, various studies have explored different key aspects. In attempt to embed GNN, Zhang et al. [55] have explored how (FPGAs) can accelerate message-passing operations in GNNs, leading to faster inference times while maintaining model accuracy. Besides, the authors also summarizes the challenges that FPGA-based GNNs accelerators need to address. Meanwhile, Wang et al. [56] proposes a quantization-based approach to address scalability challenges in GNNs. By compressing models with low-bit quantization and controlling message propagation, it reduces computational and memory costs while maintaining accuracy. The method achieves significant speedups ($5.11\times$ for INT2, $4.70\times$ for INT4), offering a cost-effective solution for deploying GNNs in NG-IoT networks. Xie et al. [57] proposed Graph sparsification and network pruning (GASSIP), a novel framework for lightweight graph neural architecture search (GNAS) tailored for resource-constrained scenarios. By combining operation-pruned architecture search and curriculum graph sparsification, GASSIP efficiently identifies optimal lightweight GNNs with reduced model parameters and sparser graphs. Extensive experiments show that GASSIP achieves comparable or superior performance with significantly fewer resources, making it highly relevant for cost-effective GNN deployment in NG-IoT applications. In [58], Wu et al. introduced Prototype-Guided Knowledge Distillation (PGKD), a novel method for distilling high-accuracy GNNs into low-latency MLPs without relying on graph edges. By leveraging class prototypes and structure-aware distillation losses, PGKD effectively transfers graph structural knowledge from GNNs to MLPs, enabling edge-free learning.
Multi-modal data processing in NG-IoT networks: Li et al. [59] proposed GNNMR, a multi-modal recommendation framework that integrates GNNs with deep mutual learning to address modality bias. By training separate GNNs on uni-modal user-item graphs and using mutual knowledge distillation, GNNMR synchronizes latent semantic relationships across modalities, improving multi-modal embeddings. It outperforms existing methods in Top-K recommendation tasks, making it suitable for NG-IoT applications requiring efficient user-item interaction modeling. Similarly, Zhang et al. [60] introduced a GNN framework for multimodal data integration, leveraging feature, decision, and deep fusion techniques to achieve high accuracy (9898%), recall (86.986.9%), and F1 score ($0.964$). This approach demonstrates robust performance in complex scenarios, highlighting the potential of multimodal GNNs for NG-IoT applications.